OpenStudy (anonymous):

Find all integer solution (n,m) for n^4 + 2n^3 + 2n^2 + 2n + 1 = m^2 please explain each step you did

OpenStudy (anonymous):

(n^2+n+1)^2 = m^2 square root both sides n^2+n+1=m hope this helped

hero (hero):

Don't forget about negative roots and don't forget to factor

OpenStudy (anonymous):

alright thanks i see if i can figure it out

hero (hero):

n^2+n+1=m -(n^2+n+1)=m